Write an equation of the line that passes through (9,6) and is perpendicular to the line whose equation is y=-1/3(x)+7.
Answer by fractalier(6550)

Well if the line you need is perpendicular to y=-1/3(x)+7, then its slope must be the negative reciprocal of -1/3 or 3.
Now use the slope-interecept method
y - y1 = m(x - x1)
y - 6 = 3(x - 9)
y - 6 = 3x - 27
y = 3x - 21